# Measuring Chain of Thought Faithfulness by Unlearning Reasoning Steps

### v3

@@ 1 Introduction
Language models (LMs) can perform various tasks accurately and verbalize some reasoning via a so-called chain of thought (CoT) Kojima et al. (2022); Wei et al. (2022), even without specialized supervised training. CoT reasoning is emerging as a powerful technique for improving the performance of LMs in complex tasks OpenAI (2024); DeepSeek-AI et al. (2025). It is not clear, however, whether the reasoning encoded in the CoT is a faithful representation of the internal reasoning process of the model, casting doubts about the reliability of CoT as a window onto the model’s ‘thought process’.
Various works set out to explore CoT faithfulness by perturbing tokens within the CoT and observing whether the contextual corruptions affect model prediction (Lanham et al., 2023; Bentham et al., 2024; Chen et al., 2024b; Madsen et al., 2024). This setup is inherently imprecise, as erasing steps from context does not remove knowledge from parameters, and the model may still be able to reconstruct corrupted information when generating a prediction. Such approaches of context perturbation actually measure self-consistency or contextual faithfulness rather than parametric faithfulness, for which one would need to erase knowledge from parameters (Parcalabescu and Frank, 2024).
@@
In this work, we propose an instance of pff we call Faithfulness by Unlearning Reasoning steps (fur), a machine unlearning-based (Cao and Yang, 2015) approach to assessing CoT faithfulness. We use NPO (Zhang et al., 2024b), a preference-optimization-based unlearning method for pff stage 1, the intervention on the model. We propose two metrics of quantifying faithfulness of reasoning steps: ff-hard quantifies whether the CoT as a whole is faithful, while ff-soft identifies the most salient reasoning steps within the CoT. Concretely, we (a) generate a CoT, (b) segment it into steps, (c) independently unlearn knowledge encoded within each step from model parameters and (d) measure the effect of erased knowledge on the models’ prediction (Figure˜1). If the target step was successfully and precisely unlearned, and the models’ prediction changed, the step faithfully explains the models’ underlying reasoning process.
@@ 2 Background and Related Work
When CoT prompted, models exhibit better performance on complex multi-hop and arithmetic reasoning tasks (Zhou et al., 2023; Fu et al., 2023b; Sprague et al., 2025) compared to being prompted directly (no-CoT). Chains of thought can be used as additional context where models can store results of intermediate hops, but they also provide additional compute irrespective of content Pfau et al. (2024); Biran et al. (2024). Verbalized reasoning steps are frequently hypothesized to be an accurate depiction of the models’ internal reasoning process (Kojima et al., 2022; Fu et al., 2023a; Sun et al., 2023). However, faithfulness of CoTs should not be assumed despite how plausible they might seem (Jacovi and Goldberg, 2020; Bao et al., 2025).
@@
Natural language explanations such as CoTs exhibit a number of issues. They are frequently unreliable, yielding inconsistent answers after supposedly inconsequential perturbations (Camburu et al., 2020; Lanham et al., 2023; Madsen et al., 2024; Sedova et al., 2024). CoTs have been shown to not align with generated answers Bao et al. (2025), they are often not useful to humans (Joshi et al., 2023) and can contain factually incorrect or hallucinated information (Kim et al., 2021, 2023; Zheng et al., 2023b; Peng et al., 2023; Zhang et al., 2024a). Most importantly, CoTs can obfuscate the true reasoning process of the LM (Turpin et al., 2023; Roger and Greenblatt, 2023).
@@
Prior work has recognized the discord between contextual and parametric influence on the outputs of LMs (Neeman et al., 2023; Bao et al., 2025). Prompting models with hypothetical or factually incorrect information causes them to change their otherwise consistently correct predictions (Kim et al., 2021, 2023; Simhi et al., 2024; Minder et al., 2025), highlighting their high sensitivity to context tokens and confounding any conclusions drawn from contextual perturbations applied to reasoning steps. The main issue with work investigating self-consistency is the possibility of the LM reconstructing information obfuscated by the contextual perturbation—despite the verbalized knowledge missing, this reasoning could still be retrieved from the latent space (Yang et al., 2024; Deng et al., 2024). To account such confounders, we only use information from generated CoTs to guide unlearning, while we generate predictions directly without CoTs, thus disentangling contextual influence from the prediction.
@@
Various tests and metrics for quantifying faithfulness of free-text explanations in LMs have previously been proposed (Lanham et al., 2023; Bentham et al., 2024; Atanasova et al., 2023; Siegel et al., 2024). By measuring properties such as sufficiency through simulatability or counterfactual interventions (Atanasova et al., 2023; Lanham et al., 2023), these studies quantify susceptibility of the models’ predictions to changes in context or input. Such approaches are valid only if there is no direct causal link between the input and prediction that bypasses the explanation, which is rare in LMs (Bao et al., 2025). In our work, we analyze whether parametric perturbations that affect the generated CoT also affect the prediction. The closest to ours are the contemporaneous works of Yeo et al. (2024) who use activation patching to measure causal effect of corrupting certain hidden states, and Zaman and Srivastava (2025) who use knowledge editing to evaluate existing (un)faithfulness metrics.
@@
Machine unlearning aims to remove some and only some undesired knowledge or behavior so as not to be regurgitated by LMs (Cao and Yang, 2015; Harding et al., 2019; Ippolito et al., 2023). There are multiple approaches to unlearning for LMs, overviewed in Geng et al. (2025) and Appendix A. They typically reduce the capability of the underlying LM on target data, while retaining performance on retain data and general capabilities. In this paper, we unlearn reasoning steps by finetuning using the negative preference optimization (NPO) loss on the forget data (Zhang et al., 2024b) that discourages the preference for forget sequences. We add it to the KL divergence between the original and “unlearned” model’s predictions on the retain set Chen and Yang (2023); Yao et al. (2024). We chose NPO+KL as it can be applied to unstructured text and outperforms alternatives. More details in §4.1.
@@
A line of work has analyzed the sensitivity of models to perturbations applied to reasoning steps (Lanham et al., 2023; Bentham et al., 2024; Chen et al., 2024b; Madsen et al., 2024, inter alia) under the guise of faithfulness. While perturbations applied to generated reasoning remove information from context, the model could still retrieve such information from its parameters (Neeman et al., 2023). Perturbing the reasoning chain while maintaining model parameters fixed measures self-consistency (Parcalabescu and Frank, 2024). Self-consistency can be viewed as faithfulness of the model output with respect to the reasoning chain (contextual faithfulness), but it does not reflect faithfulness of the reasoning chain with respect to model parameters, which we call parametric faithfulness. Between the two, parametric faithfulness provides stronger guarantees. Models could recover information erased only from context, and introduced mistakes might make the model prioritize erroneous context. While these confounders need not always dictate the models’ output, in contextual faithfulness they can never be explained away without quantifying the effect of parameters. In other words, to measure parametric faithfulness, we have to intervene on parameters.
@@ 4.1 Parameter Intervention
We erase knowledge contained in the verbalized reasoning steps using a preference-optimization unlearning method, NPO (Zhang et al., 2024b). Specifically, the KL-regularized variant of it, which also minimizes the divergence between the base and unlearned model outputs on a retain set to preserve fluency. [^2] We experimented with NPO+grad-diff, but results were always slightly worse than NPO+KL. We explored ROME and MEMIT (Meng et al., 2022, 2023), but they require a structured format, and do not perform well under paraphrases. We provide a detailed overview in Appendix˜A for readers unfamiliar with NPO+KL.
@@
We unlearn each reasoning step individually, for a total of $5$ iterations, and refer to the model obtained after unlearning the $i$-th reasoning step alone as $\mathcal{M}^{(i)^{*}}$. One unlearning iteration refers to a pass over $\mathcal{D}_{\text{FG}}$. We only update the second FF2 matrix of the Transformer MLPs, as this layer was found to act as a memory store (Geva et al., 2021b; Meng et al., 2022) and model editing methods frequently target it to update information (Meng et al., 2022, 2023; Hong et al., 2024). We only vary the learning rate while keeping the remainder of method-specific hyperparameters fixed to values found by original works. We report them in Appendix˜E.
@@ 4.2 Controls
Unlearning is deemed successful if the target information is removed (high efficacy), but the model retains its general capabilities, fluency, and performance on non-forgotten in-domain data (high specificity) (Gandikota et al., 2024). We adapt these criteria for unlearning methods within fur.
@@
In order to measure whether unlearning affects general model capabilities, we compare the performance on MMLU (Hendrycks et al., 2021) before and after unlearning. Due to prohibitive costs of evaluating few-shot MMLU for each instance and unlearned CoT step, we (1) opt for zero-shot evaluation as the instruction-tuned models report good performance in this setup, and (2) MMLU score of the model after unlearning each step from $10$ randomly selected CoTs ($\approx 50$ unlearning steps).
@@ 4.3 Faithfulness Measurements
Perfectly determining whether a reasoning chain constitutes a faithful explanation is difficult. Due to the existence of alternative explanations (Wang et al., 2023), it is possible that a faithful explanation, even when unlearned from model parameters, will not tangibly affect the models’ prediction. Therefore, we do not expect ff-hard to have perfect recall. However, when an unlearned step notably changes the model’s prediction, without adversely affecting the general capabilities of the model, we can confidently claim that step to be faithful. For the remaining $100$-$f$-hard instances, there are three possibilities: (1) fur failed to uncover and unlearn the true reasoning path, (2) the model used multiple valid reasoning paths, and unlearning one did not significantly affect its prediction, or (3) the model was genuinely unfaithful in its explanation. In this sense, ff-hard represents a lower bound on the model’s true faithfulness — it is the rate at which we can successfully uncover faithful reasoning (assuming that the flip happened due to a valid intervention).
@@ 5 Experimental Setup
We use four representative instruction-tuned models from three families: LLaMA-3-8B-Instruct and Llama-3.2-3B-Instruct Touvron et al. (2023), Mistral-7B-Instruct-v0.2 Jiang et al. (2023), and Phi-3-mini-4k-Instruct Abdin et al. (2024).
@@
We employ five diverse multi-hop datasets: OpenbookQA (Book; Mihaylov et al., 2018), ARC-Challenge (Arc-ch; Clark et al., 2018), StrategyQA (SQA; Geva et al., 2021a), TruthfulQA (TQA; Lin et al., 2022) and the Sports understanding subtask of BigBench-Hard Srivastava et al. (2023). These datasets span a variety of domains, necessitating knowledge of science, sports, geography, health, law, finance and logic. We choose MCQA as the target task as it simplifies analysis of how the models’ predictive distribution shifts after unlearning due to availability of alternative answers. To retain comparable sizes, and due to expensive runtime of unlearning each CoT step, we select a subset of $250$ instances from the test split of each dataset to balance the question sources. [^6] For SQA, we use instances from the validation split due to the availability of labels. Sports has a total of $248$ instances. Details of datasets and models are in Appendix˜C.
@@
We use a two-step prompting approach Bowman et al. (2022); Lanham et al. (2023), where the model is first prompted to generate the CoT based on the question and answer options, and subsequently prompted to complete the answer letter based on the question, answer choices, and the CoT. We use greedy decoding when generating, producing a single CoT for each instance. For the prompts used, see Appendix˜D.
@@
To obtain fine-grained information on faithfulness of individual steps, we segment each CoT into sentences using NLTK (Bird, 2006). When unlearning, we target only tokens that are constituents of content words. [^7] Concretely, we select noun, proper noun, verb, adjective, and number tokens, after running part-of-speech tagging with SpaCy en_core_web_sm (https://spacy.io/). We opt for this approach so as to not unlearn the capability to verbalize reasoning from the models, but only knowledge within the steps, which we frequently observed prior to making this modification.
@@ 6.2 Does Unlearning Change Predictions?
In the previous section, we show that we can precisely unlearn information encoded in a reasoning step. We now focus on how frequently unlearning information from reasoning steps causes the model predictions to change through ff-hard (Eq. 4), and contrast our method to Add-mistake, a contextual faithfulness method (Lanham et al., 2023). Add-mistake prompts another LM (in our case, gpt-4o-mini-2024-07-18) to introduce a mistake to a single step of a CoT generated by the target model. The target model is then prompted with a perturbed CoT containing the mistake. If the prediction of the model changes, the CoT is considered faithful. We report the results of instance-level faithfulness for fur and Add-mistake in Table˜2.888We explore and comment other baselines in Appendix B. We compare to this popular baseline, as ground-truth faithfulness is not available and cannot be directly labeled since the model’s internal process is unknown. Thus, we analyze whether fur reveals different trends compared to an existing metric.
@@ 6.3 Does Unlearning Change Reasoning?
To quantitatively assess how frequently the verbalized reasoning of the model changes post-unlearning, we employ an LLM-as-a-judge (Zheng et al., 2023a) to verify if unlearning caused the generated CoT to support a different answer, indicating deeper unlearning, or if the change in model prediction is not reflected in reasoning Cohen et al. (2024). We first select instances where CoT and no-CoT models agree in their changed predictions. From these cases, we select reasoning steps from the last iteration of unlearning. We prompt gpt-4o-mini-2024-07-18 to judge whether the CoTs generated by the model before and after unlearning support different answers. We report results in Table˜3 and detail our setup in Appendix˜H.

### v4

@@ 1 Introduction
Language models (LMs) can perform various tasks accurately and verbalize some reasoning via a so-called chain of thought (CoT) kojima2022large; wei2022chain, even without specialized supervised training. CoT reasoning is emerging as a powerful technique for improving the performance of LMs in complex tasks o1; deepseek2025r1. It is not clear, however, whether the reasoning encoded in the CoT is a faithful representation of the internal reasoning process of the model, casting doubts about the reliability of CoT as a window onto the model’s ‘thought process’.
Various works set out to explore CoT faithfulness by perturbing tokens within the CoT and observing whether the contextual corruptions affect model prediction (lanham2023measuring; DBLP:journals/tmlr/BenthamSM24; chen2024counterfactual; madsen2024self). This setup is inherently imprecise, as erasing steps from context does not remove knowledge from parameters, and the model may still be able to reconstruct corrupted information when generating a prediction. Such approaches of context perturbation actually measure self-consistency or contextual faithfulness rather than parametric faithfulness, for which one would need to erase knowledge from parameters (parcalabescu2023measuring).
@@
In this work, we propose an instance of pff we call Faithfulness by Unlearning Reasoning steps (fur), a machine unlearning-based (cao2015forget) approach to assessing CoT faithfulness. We use NPO (zhang2024npo), a preference-optimization-based unlearning method for pff stage 1, the intervention on the model. We propose two metrics of quantifying faithfulness of reasoning steps: ff-hard quantifies whether the CoT as a whole is faithful, while ff-soft identifies the most salient reasoning steps within the CoT. Concretely, we (a) generate a CoT, (b) segment it into steps, (c) independently unlearn knowledge encoded within each step from model parameters and (d) measure the effect of erased knowledge on the models’ prediction (Figure˜1). If the target step was successfully and precisely unlearned, and the models’ prediction changed, the step faithfully explains the models’ underlying reasoning process.
@@ 2 Background and Related Work
When CoT prompted, models exhibit better performance on complex multi-hop and arithmetic reasoning tasks (zhou2023complex; fu2023complexity; sprague2024cotmath) compared to being prompted directly (no-CoT). Chains of thought can be used as additional context where models can store results of intermediate hops, but they also provide additional compute irrespective of content pfau2024hidden; biran2024hopping. Verbalized reasoning steps are frequently hypothesized to be an accurate depiction of the models’ internal reasoning process (kojima2022large; fu2023specializing; sun2023recitation). However, faithfulness of CoTs should not be assumed despite how plausible they might seem (jacovi-goldberg-2020-towards; bao2024llms).
@@
Natural language explanations such as CoTs exhibit a number of issues. They are frequently unreliable, yielding inconsistent answers after supposedly inconsequential perturbations (camburu2020adversarial; lanham2023measuring; madsen2024self; sedova2024ambiguity). CoTs have been shown to not align with generated answers bao2024llms, they are often not useful to humans (joshi2023useful) and can contain factually incorrect or hallucinated information (kim2021presupposition; kim2023questionable; zheng2023does; peng2302check; zhang2024snowball). Most importantly, CoTs can obfuscate the true reasoning process of the LM (turpin2023unfaithful; roger2023preventing).
@@
Prior work has recognized the discord between contextual and parametric influence on the outputs of LMs (neeman-etal-2023-disentqa; bao2024llms). Prompting models with hypothetical or factually incorrect information causes them to change their otherwise consistently correct predictions (kim2021presupposition; kim2023questionable; simhi2024distinguishing; minder2025controllable), highlighting their high sensitivity to context tokens and confounding any conclusions drawn from contextual perturbations applied to reasoning steps. The main issue with work investigating self-consistency is the possibility of the LM reconstructing information obfuscated by the contextual perturbation—despite the verbalized knowledge missing, this reasoning could still be retrieved from the latent space (yang2024latently; deng2024explicit). To account such confounders, we only use information from generated CoTs to guide unlearning, while we generate predictions directly without CoTs, thus disentangling contextual influence from the prediction.
@@
Various tests and metrics for quantifying faithfulness of free-text explanations in LMs have previously been proposed (lanham2023measuring; DBLP:journals/tmlr/BenthamSM24; atanasova2023faithfulness; siegel2024probabilities). By measuring properties such as sufficiency through simulatability or counterfactual interventions (atanasova2023faithfulness; lanham2023measuring), these studies quantify susceptibility of the models’ predictions to changes in context or input. Such approaches are valid only if there is no direct causal link between the input and prediction that bypasses the explanation, which is rare in LMs (bao2024llms). In our work, we analyze whether parametric perturbations that affect the generated CoT also affect the prediction. The closest to ours are the contemporaneous works of yao2024patching who use activation patching to measure causal effect of corrupting certain hidden states, and zaman2025causal who use knowledge editing to evaluate existing (un)faithfulness metrics.
@@
Machine unlearning aims to remove some and only some undesired knowledge or behavior so as not to be regurgitated by LMs (cao2015forget; harding2019understanding; ippolito2023preventing). There are multiple approaches to unlearning for LMs, overviewed in geng2025comprehensivesurveymachineunlearning and Appendix A. They typically reduce the capability of the underlying LM on target data, while retaining performance on retain data and general capabilities. In this paper, we unlearn reasoning steps by finetuning using the negative preference optimization (NPO) loss on the forget data (zhang2024npo) that discourages the preference for forget sequences. We add it to the KL divergence between the original and “unlearned” model’s predictions on the retain set chen-yang-2023-unlearn; DBLP:conf/nips/YaoXL24. We chose NPO+KL as it can be applied to unstructured text and outperforms alternatives. More details in §4.1.
@@
A line of work has analyzed the sensitivity of models to perturbations applied to reasoning steps (lanham2023measuring; DBLP:journals/tmlr/BenthamSM24; chen2024counterfactual; madsen2024self, inter alia) under the guise of faithfulness. While perturbations applied to generated reasoning remove information from context, the model could still retrieve such information from its parameters (neeman-etal-2023-disentqa). Perturbing the reasoning chain while maintaining model parameters fixed measures self-consistency (parcalabescu2023measuring). Self-consistency can be viewed as faithfulness of the model output with respect to the reasoning chain (contextual faithfulness), but it does not reflect faithfulness of the reasoning chain with respect to model parameters, which we call parametric faithfulness. Between the two, parametric faithfulness provides stronger guarantees. Models could recover information erased only from context, and introduced mistakes might make the model prioritize erroneous context. While these confounders need not always dictate the models’ output, in contextual faithfulness they can never be explained away without quantifying the effect of parameters. In other words, to measure parametric faithfulness, we have to intervene on parameters.
@@ 4.1 Parameter Intervention
We erase knowledge contained in the verbalized reasoning steps using a preference-optimization unlearning method, NPO (zhang2024npo). Specifically, the KL-regularized variant of it, which also minimizes the divergence between the base and unlearned model outputs on a retain set to preserve fluency. [^2] We experimented with NPO+grad-diff, but results were always slightly worse than NPO+KL. We explored ROME and MEMIT (meng2022locating; meng2023memit), but they require a structured format, and do not perform well under paraphrases. We provide a detailed overview in Appendix˜A for readers unfamiliar with NPO+KL.
@@
We unlearn each reasoning step individually, for a total of $5$ iterations, and refer to the model obtained after unlearning the $i$-th reasoning step alone as $\mathcal{M}^{(i)^{*}}$. One unlearning iteration refers to a pass over $\mathcal{D}_{\text{FG}}$. We only update the second FF2 matrix of the Transformer MLPs, as this layer was found to act as a memory store (geva2020transformer; meng2022locating) and model editing methods frequently target it to update information (meng2022locating; meng2023memit; hong2024intrinsic). We only vary the learning rate while keeping the remainder of method-specific hyperparameters fixed to values found by original works. We report them in Appendix˜E.
@@ 4.2 Controls
Unlearning is deemed successful if the target information is removed (high efficacy), but the model retains its general capabilities, fluency, and performance on non-forgotten in-domain data (high specificity) (gandikota2024erasing). We adapt these criteria for unlearning methods within fur.
@@
In order to measure whether unlearning affects general model capabilities, we compare the performance on MMLU (hendrycks2020measuring) before and after unlearning. Due to prohibitive costs of evaluating few-shot MMLU for each instance and unlearned CoT step, we (1) opt for zero-shot evaluation as the instruction-tuned models report good performance in this setup, and (2) MMLU score of the model after unlearning each step from $10$ randomly selected CoTs ($\approx 50$ unlearning steps).
@@ 4.3 Faithfulness Measurements
Perfectly determining whether a reasoning chain constitutes a faithful explanation is difficult. Due to the existence of alternative explanations (wang2022self), it is possible that a faithful explanation, even when unlearned from model parameters, will not tangibly affect the models’ prediction. Therefore, we do not expect ff-hard to have perfect recall. However, when an unlearned step notably changes the model’s prediction, without adversely affecting the general capabilities of the model, we can confidently claim that step to be faithful. For the remaining $100$-$f$-hard instances, there are three possibilities: (1) fur failed to uncover and unlearn the true reasoning path, (2) the model used multiple valid reasoning paths, and unlearning one did not significantly affect its prediction, or (3) the model was genuinely unfaithful in its explanation. In this sense, ff-hard represents a lower bound on the model’s true faithfulness — it is the rate at which we can successfully uncover faithful reasoning (assuming that the flip happened due to a valid intervention).
@@ 5 Experimental Setup
We use four representative instruction-tuned models from three families: LLaMA-3-8B-Instruct and Llama-3.2-3B-Instruct touvron2023llama, Mistral-7B-Instruct-v0.2 jiang2023mistral, and Phi-3-mini-4k-Instruct DBLP:journals/corr/abs-2404-14219.
@@
We employ five diverse multi-hop datasets: OpenbookQA (Book; mihaylov-etal-2018-suit), ARC-Challenge (Arc-ch; DBLP:journals/corr/abs-1803-05457), StrategyQA (SQA; geva2021did), TruthfulQA (TQA; lin2022truthfulqa) and the Sports understanding subtask of BigBench-Hard srivastava2023bbh. These datasets span a variety of domains, necessitating knowledge of science, sports, geography, health, law, finance and logic. We choose MCQA as the target task as it simplifies analysis of how the models’ predictive distribution shifts after unlearning due to availability of alternative answers. To retain comparable sizes, and due to expensive runtime of unlearning each CoT step, we select a subset of $250$ instances from the test split of each dataset to balance the question sources. [^6] For SQA, we use instances from the validation split due to the availability of labels. Sports has a total of $248$ instances. Details of datasets and models are in Appendix˜C.
@@
We use a two-step prompting approach bowman2022scalable; lanham2023measuring, where the model is first prompted to generate the CoT based on the question and answer options, and subsequently prompted to complete the answer letter based on the question, answer choices, and the CoT. We use greedy decoding when generating, producing a single CoT for each instance. For the prompts used, see Appendix˜D.
@@
To obtain fine-grained information on faithfulness of individual steps, we segment each CoT into sentences using NLTK (bird2006nltk). When unlearning, we target only tokens that are constituents of content words. [^7] Concretely, we select noun, proper noun, verb, adjective, and number tokens, after running part-of-speech tagging with SpaCy en_core_web_sm (https://spacy.io/). We opt for this approach so as to not unlearn the capability to verbalize reasoning from the models, but only knowledge within the steps, which we frequently observed prior to making this modification.
@@ 6.2 Does Unlearning Change Predictions?
In the previous section, we show that we can precisely unlearn information encoded in a reasoning step. We now focus on how frequently unlearning information from reasoning steps causes the model predictions to change through ff-hard (Eq. 4), and contrast our method to Add-mistake, a contextual faithfulness method (lanham2023measuring). Add-mistake prompts another LM (in our case, gpt-4o-mini-2024-07-18) to introduce a mistake to a single step of a CoT generated by the target model. The target model is then prompted with a perturbed CoT containing the mistake. If the prediction of the model changes, the CoT is considered faithful. We report the results of instance-level faithfulness for fur and Add-mistake in Table˜2.888We explore and comment other baselines in Appendix B. We compare to this popular baseline, as ground-truth faithfulness is not available and cannot be directly labeled since the model’s internal process is unknown. Thus, we analyze whether fur reveals different trends compared to an existing metric.
@@ 6.3 Does Unlearning Change Reasoning?
To quantitatively assess how frequently the verbalized reasoning of the model changes post-unlearning, we employ an LLM-as-a-judge (zheng2023llmasjudge) to verify if unlearning caused the generated CoT to support a different answer, indicating deeper unlearning, or if the change in model prediction is not reflected in reasoning cohen2024ripple. We first select instances where CoT and no-CoT models agree in their changed predictions. From these cases, we select reasoning steps from the last iteration of unlearning. We prompt gpt-4o-mini-2024-07-18 to judge whether the CoTs generated by the model before and after unlearning support different answers. We report results in Table˜3 and detail our setup in Appendix˜H.